A compound composed of only carbon and hydrogen is
Answer #1

7)

Given :

mass is in mass percentage. Let the total mass =100 g

=> mass of Hydrogen = 7.743 g

mass of Carbon = 100 - 7.743 = 92.257 g

moles of Carbon , m1 = mass / molecular mass = 92.257 / 12 = 7.69 moles

moles of hydrogen , m2 = mass / molecular mass = 7.743 / 1 = 7.743 moles

  • Now for finding the structures, we have divide the moles with the smallest moles, here mole of Hydrogen is the smallest.

=> Ratio of the carbon atoms = 7.69 / 7.743 = 0.993 = 1 (approx)

=> Ratio of the hydrogen atoms = 7.743 / 7.743 = 1

so the emperical formula will be CH.

Structural formula :

As the valency of C = 4 , it will make 4 bonds, but in the above formula it makes only one.

\therefore structural formula will be C2H2

Lewis structure is given below :

H-CEC-H

8)

If you draw a Lewis dot structure of I3- (triiodide ion, which is linear) and Br3- (tribromide) you'll see that the central I atom forms two bonds, and has three non-bonding pairs of electrons. That's a total of five electron pairs - more than an octet. A lot of compounds violate the octet rule that way, but the first long-row elements like F can't. The reason is that the lower elements like iodine can use an available d-orbital to accomodate the additional pair of electrons (sp3d hybridization), but F simply doesn't have a d-orbital that it can use.